OpenStudy (anonymous):

put in exponetial form sqrt^7sqrtx

OpenStudy (anonymous):

\[\sqrt{\sqrt[7]{x}}\]

OpenStudy (anonymous):

choices are, \[x ^{1/5}...x ^{1/14}....x ^{1/9}....x ^{1/8}\]

OpenStudy (anonymous):

second one

OpenStudy (anonymous):

\[(x^{\frac{1}{7}})^{\frac{1}{2}}\] might be a way to see it